# Biased random walk on dynamical percolation

### v2

@@ 3 Speed and Einstein relation
$=\sum_{m\in\mathbb{N}}\sum_{\begin{subarray}{c}k>\ell\\ k+\ell\leq m\end{subarray}}(k-\ell)\left(\frac{2d}{Z_{\lambda}}\right)^{m}\left(e^{\lambda(k-\ell)}-e^{-\lambda(k-\ell)}\right)\mathbb{P}_{0}((R,L)=(k,\ell),\mathcal{U}(\tau_{1})=m).$ (3.17)
Since all the terms of the sum in (3.17) are positive, we can infer a strictly positive speed using that $\mathbb{P}_{0}((R,L)=(2,1))>0$. ∎
@@
Theorem 1.2 is now an immediate consequence of Propositions 3.1 and 3.5 and Lemma 3.4. ∎
@@ 4.1 Speed for $d=1$
Let $(X,\eta)$ be a totally asymmetric biased random walk in dynamical percolation on $\mathbb{Z}$ with parameters $\mu$ and $p$ that jumps to the right at rate $1$. Then the speed $\overline{v}$ satisfies
@@
Suppose $X_{0}=0$ and $\eta_{0}\sim\pi$. Let $S$ be the first time that $X$ jumps along the edge $e=\{0,1\}$. Then $\overline{v}=\mathbb{E}\!\left[S\right]^{-1}$. To compute $\mathbb{E}\!\left[S\right]$ we describe a way to realise the evolution of the state of $e$ over time. Let $(\xi_{i})_{i\in\mathbb{N}}$ be i.i.d. Bernoulli-$p$-distributed random variables. Let $(T_{i})_{i\in\mathbb{N}}$ and $(S_{j})_{j\in\mathbb{N}}$ denote the rate $\mu$ and rate $1$ Poisson clocks that determine when the edge $e$ is updated and when the random walk attempts a move to the right, respectively. Taking $T_{0}:=0$, we assign to each interval $[T_{i-1},T_{i}]$ the label $\xi_{i}$. If $\xi_{i}=1$, the edge $e$ is open and otherwise it is closed. We let
$Z:=\inf\left\{j\in\mathbb{N}\colon S_{j}\in[T_{i-1},T_{i}]\text{ for some }i\in\mathbb{N}\text{ with }\xi_{i}=1\right\}.$ (4.4)
In particular, we have $S=S_{Z}$. By the memoryless property of the exponential distribution we get
$\mathbb{P}(Z=j|Z>j-1)=\frac{\mu p}{\mu+1}\quad\text{ and }\quad\mathbb{P}(Z=1)=p.$ (4.5)
Therefore, for all $j>1$, we have that
$\mathbb{P}(Z=j)$ $=\mathbb{P}(Z=j|Z>j-1)(1-p)\prod_{i=2}^{j-1}\mathbb{P}(Z>i|Z>i-1)$
$=(1-p)\Big{(}1-\frac{\mu p}{\mu+1}\Big{)}^{j-2}\frac{\mu p}{\mu+1}.$
It then follows that
$\mathbb{E}\!\left[Z\right]=\frac{1-p+\mu}{\mu p}.$ (4.6)
Let $(\chi_{i})_{i\in\mathbb{N}}$ with $\chi_{i}=S_{i+1}-S_{i}$ denote the inter-arrival times of the Poisson process $(S_{j})_{j\in\mathbb{N}}$. In particular, $(\chi_{i})_{i\in\mathbb{N}}$ are i.i.d. Exponential-$1$-distributed random variables. Observe that $Z$ is a stopping time with respect to $(\chi_{i})_{i\in\mathbb{N}}$, when we consider the enlarged filtration which contains also the process $(T_{i})_{i\in\mathbb{N}}$ which is independent of $(\chi_{i})_{i\in\mathbb{N}}$. Therefore, applying Wald’s identity and using (4.6), we conclude
$\mathbb{E}[S]=\mathbb{E}[\chi_{1}]\mathbb{E}[Z]=\frac{1-p+\mu}{\mu p}$
and this finishes the proof. ∎
@@
Otherwise, we add to $\overline{I}_{t}$ the copy $e_{i,j}$ of $e_{i}$ with the smallest index $j$ such that $e_{i,j}\notin\overline{I}_{t_{-}}$. If the random walks are at the same position, but examine different edges, we add for both copies of the edges with the smallest index as above to the modified infected set. When the two random walks are at different positions, recall that according to the coupling, the two random walks perform jumps according to independent exponential $1$ clocks. Whenever an edge is examined by one of the two random walks, we add the respective copy of it to the modified infected set.
@@
Next, we investigate the speed when $\lambda\rightarrow\infty$. Since the $\mathbf{P}$-probability for $(X_{t})$ to attempt a jump in the $-\textup{e}_{1}$ direction until time $\overline{\tau}$ goes to 00 as $\lambda_{1}\rightarrow\infty$, and the speed is increasing in $\lambda$,
@@
where $\overline{v}$ denotes the speed of the totally biased random walk, i.e.  whenever the clock associated to the random walker at $x$ rings, it attempts a jump to $x+1$. Lemma 4.4 concludes the proof. ∎
@@ 4.2 Asymptotic expression for the speed
In the following, with a slight abuse of notation, we identify $\mathbb{Z}$ with the $\textup{e}_{1}$-axis of $\mathbb{Z}^{d}$, and for $x\in\mathbb{Z}$, we write $x+\textup{e}_{1}$ for the edge $(x,x+1)$. Recall that $Z_{\lambda}=e^{\lambda}+e^{-\lambda}+2d-2$.
The purpose of this coupling is to allow us to transfer the dynamics on $\mathbb{Z}^{d}$ to a one-dimensional system, which we can study explicitly. More precisely, we consider three Poisson processes $\mathcal{P}^{1},\mathcal{P}^{2},\mathcal{P}^{3}$. The processes $\mathcal{P}^{1}$ and $\mathcal{P}^{2}$ correspond to jumps in the ±$\pm\textup{e}_{1}$ direction, while $\mathcal{P}^{3}$ corresponds to all other directions of $\mathbb{Z}^{d}$. At the points of $\mathcal{P}^{1}$ and $\mathcal{P}^{2}$, we let the two processes $X$ and $Y$ evolve together, and we update their respective infected sets together. We stop the coupling either at the first jump in direction $-\textup{e}_{1}$ or the second time a point in $\mathcal{P}^{3}$ occurs, and continue the processes afterwards according to the correct marginal transition rates; see below. Our key observation is that as long as the edge to the right of the walker’s current location has not been examined before, we can identify the state of the edge on $\mathbb{Z}$ with the same $\textup{e}_{1}$-coordinate. We make this rigorous in the last part of the definition below.
@@
Let $\widetilde{\mu}=\mu+p(2d-2)Z_{\lambda}^{-1}$. Both $X$ and  $Y$ start from 00. We let the environment $\eta$ evolve according to dynamical percolation on $\mathbb{Z}^{d}$ with parameters $\mu,p$ and $\eta_{0}\sim\pi_{p}$. The edges to the left of 00 in the environment $\xi$ update according to dynamical percolation with parameters $\widetilde{\mu},p$.
@@
$\mathbb{E}\!\left[{\text{\Large$\mathfrak{1}$}}(A)\cdot\mathcal{U}_{a}(\tau_{1})\right]\leq\mathbb{E}\!\left[\mathcal{U}_{a}(\tau_{1})\left(1-\left(1-e^{-\lambda}Z_{\lambda}^{-1}\right)^{\mathcal{U}_{a}(\tau_{1})}\right)\right]$
$+\mathbb{E}\!\left[\mathcal{U}_{a}(\tau_{1})\left(1-\left(1-(2d-2)Z_{\lambda}^{-1}\right)^{\mathcal{U}_{a}(\tau_{1})}-\mathcal{U}_{a}(\tau_{1})(2d-2)Z_{\lambda}^{-1}(1-(2d-2)Z_{\lambda}^{-1})^{\mathcal{U}_{a}(\tau_{1})-1}\right)\right].$
Using that for all $x\in(0,1)$ and $a\in\mathbb{N}$ we have $(1-x)^{a}\geq 1-ax$, we get
$\mathbb{E}\!\left[{\text{\Large$\mathfrak{1}$}}(A)\cdot\mathcal{U}_{a}(\tau_{1})\right]\leq\mathbb{E}\!\left[(\mathcal{U}_{a}(\tau_{1}))^{2}e^{-\lambda}Z_{\lambda}^{-1}+\mathcal{U}_{a}(\tau_{1})(\mathcal{U}_{a}(\tau_{1})-1)(2d-2)^{2}Z_{\lambda}^{-2}\right].$
Since $Z_{\lambda}^{-1}=\mathcal{O}(e^{-\lambda})$ and by Lemma 2.3 we have for a positive constant $C_{\mu}$ that $\mathbb{E}\!\left[(\mathcal{U}_{a}(\tau_{1}))^{2}\right]\leq C_{\mu}<\infty$, it follows that
@@
We now want to compare the speed of $Z$ to the speed of $\overline{Y}$. We couple $Z$ and $\overline{Y}$ by letting them evolve together until the first time that $\overline{Y}$ attempts a jump to the left. Afterwards, they both attempt jumps at the same times and they use the same Poisson process to remove copies of edges from their infected sets. We let $\tau_{1}$ be their first regeneration time. Let $A$ be the event that the walker $\overline{Y}$ attempts at least one jump to the left before time $\tau_{1}$ and let $\mathcal{U}_{a}(\tau_{1})$ be the total number of jump attempts before time $\tau_{1}$. Therefore we deduce
$\left|\frac{\overline{\mu}p}{1-p+\overline{\mu}}-{v}^{\overline{Y}}(\lambda)\right|$ $\leq\frac{1}{\mathbb{E}\!\left[\tau_{1}\right]}\mathbb{E}\!\left[|Z_{\tau_{1}}-\overline{Y}_{\tau_{1}}|{\text{\Large$\mathfrak{1}$}}(A)\right]\leq\frac{2}{\mathbb{E}\!\left[\tau_{1}\right]}\mathbb{E}\!\left[\mathcal{U}_{a}(\tau_{1}){\text{\Large$\mathfrak{1}$}}(A)\right]$
$\leq\frac{2}{\mathbb{E}\!\left[\tau_{1}\right]}\mathbb{E}\!\left[(\mathcal{U}_{a}(\tau_{1}))^{2}\cdot\frac{e^{-\lambda}}{e^{\lambda}+e^{-\lambda}}\right]\leq C^{\prime}\exp(-2\lambda),$
where $C^{\prime}$ is a constant only depending on $\mu$ and where for the third inequality we used a union bound and for the last one we used that $\mathcal{U}_{a}(\tau_{1})$ has exponential tails by Lemma 2.3.

### v3

@@ 3 Speed and Einstein relation
$\begin{split}&\quad\mathbb{E}_{\lambda}\!\left[X_{t}^{1}\right]=\mathbb{E}_{0}\left[\left(R(t)-L(t)\right)e^{\lambda(R_{a}(t)-L_{a}(t))}\left(\frac{2d}{Z_{\lambda}}\right)^{\mathcal{U}_{a}(t)}\right]\\ &=\mathbb{E}_{0}\left[\left(R(t)-L(t)\right)e^{\lambda(R(t)+R_{\rm supp}(t)-L(t)-L_{\rm supp}(t))}\left(\frac{2d}{Z_{\lambda}}\right)^{\mathcal{U}_{a}(t)}\right]\\ &=\mathbb{E}_{0}\left[\left(L(t)-R(t)\right)e^{\lambda(L(t)+R_{\rm supp}(t)-R(t)-L_{\rm supp}(t))}\left(\frac{2d}{Z_{\lambda}}\right)^{\mathcal{U}_{a}(t)}\right]\\ &=\frac{1}{2}\mathbb{E}_{0}\left[\left(R(t)-L(t)\right)e^{\lambda(R_{\rm supp}(t)-L_{\rm supp}(t))}\left(e^{\lambda(R(t)-L(t))}-e^{-\lambda(R(t)-L(t))}\right)\left(\frac{2d}{Z_{\lambda}}\right)^{\mathcal{U}_{a}(t)}\right]\\ &>0,\end{split}$ (3.24)
where the last inequality follows since the function $x\mapsto x\left(e^{\lambda x}-e^{-\lambda x}\right)$ is non-negative and strictly positive for $x\neq 0$ and we have $\mathbb{P}_{0}(R(t)\neq L(t))>0$. (The fourth equality comes from adding the two previous lines and dividing by $2$). Hence, we know that for fixed $t$, $\mathbb{E}_{\lambda}\!\left[X_{t}^{1}\right]>0$ and we have to show that $\mathbb{E}_{\lambda}\!\left[X_{\tau_{1}}^{1}\right]>0$. Going back to (3.24),
$\quad\mathbb{E}_{\lambda}\!\left[X_{t}^{1}\right]$
$=\frac{1}{2}\mathbb{E}_{0}\left[\left(R(t)-L(t)\right)e^{\lambda(R_{\rm supp}(t)-L_{\rm supp}(t))}\left(e^{\lambda(R(t)-L(t))}-e^{-\lambda(R(t)-L(t))}\right)\left(\frac{2d}{Z_{\lambda}}\right)^{\mathcal{U}_{a}(t)}\right]$
$=\frac{1}{2}\mathbb{E}_{\lambda}\left[\left(R(t)-L(t)\right)e^{\lambda(R_{\rm supp}(t)-L_{\rm supp}(t))}\left(e^{\lambda(R(t)-L(t))}-e^{-\lambda(R(t)-L(t))}\right)e^{-\lambda(R_{a}(t)-L_{a}(t))}\right]\,,$
where we used (LABEL:prodRN1) for the last equality. Hence
$\quad\mathbb{E}_{\lambda}\!\left[X_{t}^{1}\right]$
$=\frac{1}{2}\mathbb{E}_{\lambda}\left[\left(R(t)-L(t)\right)e^{-\lambda(R(t)-L(t))}\left(e^{\lambda(R(t)-L(t))}-e^{-\lambda(R(t)-L(t))}\right)\right]$
$=\frac{1}{2}\mathbb{E}_{\lambda}\left[\left(R(t)-L(t)\right)\left(1-e^{-2\lambda(R(t)-L(t))}\right)\right]\,.$
This implies
$\mathbb{E}_{\lambda}\!\left[X_{t}^{1}\right]=-\mathbb{E}_{\lambda}\left[\left(R(t)-L(t)\right)e^{-2\lambda(R(t)-L(t))}\right]\,.$ (3.25)
We will show that, in the same spirit as in (3.2),
$\lim\limits_{t\to\infty}\frac{1}{t}\mathbb{E}_{\lambda}\left[\left(R(t)-L(t)\right)e^{-2\lambda(R(t)-L(t))}\right]=\frac{1}{\mathbb{E}[\tau_{1}]}\mathbb{E}_{\lambda}\left[\left(R-L\right)e^{-2\lambda(R-L)}\right]$ (3.26)
Assuming (3.26), dividing by $t$ in (3.25) and letting $t\to\infty$ gives
$\mathbb{E}_{\lambda}\!\left[X_{\tau_{1}}^{1}\right]=-\mathbb{E}_{\lambda}\left[\left(R-L\right)e^{-2\lambda(R-L)}\right]=\frac{1}{2}\mathbb{E}_{\lambda}\left[\left(R-L\right)\left(1-e^{-2\lambda(R-L)}\right)\right]>0,$
since the function $x\mapsto x\left(1-e^{-2\lambda x}\right)$ is non-negative and strictly positive for $x\neq 0$ and we have $\mathbb{P}_{\lambda}(R\neq L)>0$. It remains to prove (3.26). We claim that for all $t>s$ we have
$\mathbb{E}_{\lambda}\!\left[e^{-2\lambda(R(t)-L(t))}\;\middle|\;{\cal F}_{s}\right]=e^{-2\lambda(R(s)-L(s))}\,$ (3.27)
and defer the proof. Hence, $e^{-2\lambda(R(t)-L(t))}$ can act as a Radon-Nikodym derivative. More precisely, define the probability measure ¯$\overline{\mathbb{P}}_{\lambda}$ by
$\left.\frac{d\overline{\mathbb{P}}_{\lambda}}{d\mathbb{P}_{\lambda}}\right|_{{\cal F}_{t}}=e^{-2\lambda(R(t)-L(t))}\,.$
Note that, as in the proof of (LABEL:prodRNtau), for every $i$, the regeneration time $\tau_{i}$ is a stopping time with respect to the filtration $({\cal F}_{t})_{t\geq 0}$, and ¯$\overline{\mathbb{P}}_{\lambda}$ is absolutely continuous with respect to $\mathbb{P}_{\lambda}$ on ${\cal F}_{\tau_{i}}$ (recall the beginning of the proof of Lemma 3.4). Hence
$\left.\frac{d\overline{\mathbb{P}}_{\lambda}}{d\mathbb{P}_{\lambda}}\right|_{{\cal F}_{\tau_{i}}}=e^{-2\lambda(R(\tau_{i})-L(\tau_{i}))}\,.$ (3.28)
Therefore, taking $i=1$ we get
$\mathbb{E}_{\lambda}\left[\left(R-L\right)e^{-2\lambda(R-L)}\right]=\overline{\mathbb{E}}_{\lambda}\left[R-L\right]\,,$ (3.29)
where $\overline{\mathbb{E}}_{\lambda}$ is the expectation with respect to ¯$\overline{\mathbb{P}}_{\lambda}$. Using the fact that the $(\tau_{i})$’s are regeneration times under $\mathbb{P}_{\lambda}$ and (3.28) immediately yield that the $(\tau_{i})$’s are also regeneration times under ¯$\overline{\mathbb{P}}_{\lambda}$. We can now proceed as in the proof of (3.2), replacing the probability measure $\mathbb{P}_{\lambda}$ by ¯$\overline{\mathbb{P}}_{\lambda}$, to get
$\lim\limits_{t\to\infty}\frac{1}{t}\mathbb{E}_{\lambda}\left[\left(R(t)-L(t)\right)e^{-2\lambda(R(t)-L(t))}\right]$ $=\lim\limits_{t\to\infty}\frac{1}{t}\overline{\mathbb{E}}_{\lambda}\left[R(t)-L(t)\right]$
$=\frac{1}{\mathbb{E}[\tau_{1}]}\overline{\mathbb{E}}_{\lambda}\left[R-L\right]$
$=\frac{1}{\mathbb{E}[\tau_{1}]}\mathbb{E}_{\lambda}\left[\left(R-L\right)e^{-2\lambda(R-L)}\right]\,,$
where we used (3.29) for the last equality. Finally, we prove (3.27). Note that it suffices to establish the case $s=0$ and $(X_{0},\eta_{0})=(x,\zeta)$. Omitting the starting state from the notation we have
$\mathbb{E}_{\lambda}\left[e^{-2\lambda(R(t)-L(t))}\right]$ $=\mathbb{E}_{0}\left[e^{-2\lambda(R(t)-L(t))}e^{\lambda(R_{a}(t)-L_{a}(t))}\left(\frac{2d}{Z_{\lambda}}\right)^{\mathcal{U}_{a}(t)}\right]$
$=\mathbb{E}_{0}\left[e^{-\lambda(R(t)-L(t))}e^{\lambda(R_{\rm supp}(t)-L_{\rm supp}(t))}\left(\frac{2d}{Z_{\lambda}}\right)^{\mathcal{U}_{a}(t)}\right]$
$=\mathbb{E}_{0}\left[e^{\lambda(R(t)-L(t))}e^{\lambda(R_{\rm supp}(t)-L_{\rm supp}(t))}\left(\frac{2d}{Z_{\lambda}}\right)^{\mathcal{U}_{a}(t)}\right]$
$=\mathbb{E}_{0}\left[e^{\lambda(R_{a}(t)-L_{a}(t))}\left(\frac{2d}{Z_{\lambda}}\right)^{\mathcal{U}_{a}(t)}\right]=1\,,$
where we again used the time-reversal argument in the proof of Proposition 3.6 for the second-to-last equality. ∎
@@
Theorem 1.2 is now an immediate consequence of Proposition 3.1, Proposition 3.6, Proposition 3.7 and Lemma 3.5. ∎
@@ 4.1 Speed for $d=1$
Let $(X,\eta)$ be a totally asymmetric biased random walk in dynamical percolation on $\mathbb{Z}$ with parameters $\mu$ and $p$ that attempts a jump to the right at rate $1$. Then the speed $\overline{v}$ satisfies
@@
Suppose $X_{0}=0$ and $\eta_{0}\sim\pi$. Let $S$ be the first time that $X$ jumps along the edge $e=\{0,1\}$. Then $\overline{v}=\mathbb{E}\!\left[S\right]^{-1}$. In order to compute $\mathbb{E}[S]$, we use the following recursion,
$\mathbb{E}\!\left[S\right]=p+(1-p)\Big{(}1+\frac{1}{\mu}+\mathbb{E}\!\left[S\right]\Big{)}.$
Indeed, at the time of the first attempted jump of the walk the edge $e$ is open with probability $p$. Otherwise, the walk has to wait an exponential time with parameter $\mu$ until $e$ refreshes, and when this happens the situation is the same as at time 00. This recursion readily implies
$\mathbb{E}[S]=\frac{1-p+\mu}{\mu p},$
which gives the result. ∎
@@
Otherwise, we add to $\overline{I}_{t}$ the copy $e_{i,j}$ of $e_{i}$ with the smallest index $j$ such that $e_{i,j}\notin\overline{I}_{t_{-}}$. If the random walks are at the same position, but examine different edges, we add for both the copies of the edges with the smallest index as above to the modified infected set. When the two random walks are at different positions, recall that according to the coupling, the two random walks perform jumps according to independent exponential $1$ clocks. Whenever an edge is examined by one of the two random walks, we add its respective copy to the modified infected set.
@@
Next, we investigate the speed when $\lambda\rightarrow\infty$. Let $X$ be a $\lambda$-biased walk and let $\widetilde{X}$ be a totally biased random walk as in Lemma 4.4, i.e. it only attempts jumps to the right at rate $1$. We couple $X$ and $\widetilde{X}$ by letting them both jump at the points of a Poisson process $\mathcal{P}$ of rate $1$. At each jump attempt, we toss an independent coin and with probability $e^{\lambda}/(e^{\lambda}+e^{-\lambda})$ they both attempt a jump to the right, while with the complementary probability $X$ attempts a jump to the left while $\widetilde{X}$ attempts a jump to the right. After the first time that they move in different directions, we couple their environments by using the same rate $\mu$ Poisson process for the removal of copies of edges of their respective infected sets. In this way, both infected sets always have the same size and the two walks have the same regeneration times. Letting $\tau_{1}$ be their first regeneration time, we then have
$|v(\lambda)-\overline{v}|=\frac{1}{\mathbb{E}\!\left[{{\tau_{1}}}\right]}\cdot\left|\mathbb{E}\!\left[X_{{\tau_{1}}}\right]-\mathbb{E}\!\left[\widetilde{X}_{{\tau_{1}}}\right]\right|.$
By the description of the coupling above, it is clear that given $|\mathcal{P}(\tau_{1})|$, the number of attempted jumps to the left by $X$ has the binomial distribution with parameters $|\mathcal{P}(\tau_{1})|$ and $e^{-\lambda}/(e^{\lambda}+e^{-\lambda})$. Using the above coupling we see that $X_{\tau_{1}}$ is equal to $\widetilde{X}_{\tau_{1}}$ on the event that there is no left jump by time $\tau_{1}$. We then get by a union bound
$\left|\mathbb{E}\!\left[X_{{\tau_{1}}}\right]-\mathbb{E}\!\left[\widetilde{X}_{{\tau_{1}}}\right]\right|\leq\mathbb{E}\!\left[|\mathcal{P}(\tau_{1})|^{2}\cdot\frac{e^{-\lambda}}{e^{\lambda}+e^{-\lambda}}\right]\leq C\cdot\exp(-2\lambda),$ (4.7)
where $C$ is a positive constant using also Lemma 2.3. We can thus deduce that as $\lambda\to\infty$ the above tends to 00, and hence we conclude that
@@
where $\bar{v}$ is given in Lemma 4.4. ∎
@@ 4.2 Asymptotic expression for the speed
In the following, with a slight abuse of notation, we identify $\mathbb{Z}$ with the $\textup{e}_{1}$-axis of $\mathbb{Z}^{d}$, and for $x\in\mathbb{Z}$, we write $x+\textup{e}_{1}$ for the edge $(x,x+1)$.
The purpose of this coupling is to allow us to transfer the dynamics on $\mathbb{Z}^{d}$ to a one-dimensional system, which we can study explicitly. More precisely, we consider three Poisson processes $\mathcal{P}^{1},\mathcal{P}^{2},\mathcal{P}^{3}$. The processes $\mathcal{P}^{1}$ and $\mathcal{P}^{2}$ correspond to jumps in the ±$\pm\textup{e}_{1}$ direction, while $\mathcal{P}^{3}$ corresponds to all other directions of $\mathbb{Z}^{d}$. At the points of $\mathcal{P}^{1}$ and $\mathcal{P}^{2}$, we let the two processes $X$ and $Y$ evolve together, and we update their respective infected sets together. We stop the coupling either at the first jump in direction $-\textup{e}_{1}$ or the second time a point in $\mathcal{P}^{3}$ occurs, and continue the processes afterwards according to the correct marginal transition rates; see below. Our key observation is that as long as the edge to the right of the walker’s current location has not been examined before, we can identify its state with the state of the edge on $\mathbb{Z}$ with the same $\textup{e}_{1}$-coordinate. We make this rigorous in the last part of the definition below.
@@
Let $\widetilde{\mu}=\mu+p(2d-2)Z_{\lambda}^{-1}$, where we recall that $Z_{\lambda}=e^{\lambda}+e^{-\lambda}+2d-2$. Both $X$ and  $Y$ start from 00. We let the environment $\eta$ evolve according to dynamical percolation on $\mathbb{Z}^{d}$ with parameters $\mu,p$ and $\eta_{0}\sim\pi_{p}$. The edges to the left of 00 in the environment $\xi$ update according to dynamical percolation with parameters $\widetilde{\mu},p$.
@@
$\mathbb{E}\!\left[{\text{\Large$\mathfrak{1}$}}(A)\cdot\mathcal{U}_{a}(\tau_{1})\right]\leq\mathbb{E}\!\left[(\mathcal{U}_{a}(\tau_{1}))^{2}\cdot e^{-\lambda}Z_{\lambda}^{-1}+(\mathcal{U}_{a}(\tau_{1}))^{3}\cdot(2d-2)^{2}Z_{\lambda}^{-2}\right].$
Since $Z_{\lambda}^{-1}=\mathcal{O}(e^{-\lambda})$ and by Lemma 2.3 there exists a positive constant $C_{\mu}$ such that $\mathbb{E}\!\left[(\mathcal{U}_{a}(\tau_{1}))^{3}\right]\leq C_{\mu}<\infty$, it follows that
@@
By (4.7) we also get that for a positive constant $C^{\prime}$ we have
$\left|\frac{\overline{\mu}p}{1-p+\overline{\mu}}-{v}^{\overline{Y}}(\lambda)\right|\leq C^{\prime}\exp(-2\lambda).$